OpenStudy (anonymous):

The two lines represented by the equations below are graphed on a coordinate plane. x + 6y = 12 3(x − 2) = −y − 4 Which statement best describes the two lines? (1) The lines are parallel. (2) The lines are the same line. (3) The lines are perpendicular. (4) The lines intersect at an angle other than 90°.

OpenStudy (precal):

it would be (1) if the slopes are the same. it would be (2) is the equations are the same. it would be (3) is the slopes are opposite reciprocals (in this case 6 and -1/6) it would be (4) if they lines intersect at a point

OpenStudy (anonymous):

i know but which one is it?

OpenStudy (precal):

I am trying to help you understand the concept. Find the slopes of both lines and that will tell you if it is case (1) or (3)

OpenStudy (anonymous):

i know =_= i just needa check over my answer y=-3x+2; y=-x/6=12/6; im not sure if i put it in y equals form correct

OpenStudy (anonymous):

i got 4

OpenStudy (precal):

Yes put them both equal to y and the number in front of x is your slope

OpenStudy (precal):

tell me what your y equations are and I will tell you how to finish the problem